# Dieter van Melkebeek

### v1

@@ 1 Introduction
Many an undergraduate algorithms course features Interval Scheduling to illustrate the design paradigms of greed and dynamic programming. The unweighted version admits an appealing greedy strategy that can be shown correct via several types of arguments (charging, exchange, greedy stays ahead). The weighted version exhibits a straightforward principle of optimality and demonstrates one of the measures that can be taken to keep the number of subproblems in a dynamic program small – ordering the input components judiciously. However, the appropriate order for the memoized dynamic program differs from the order in the greedy algorithm. This tends to confuse students and gives some the impression that, even though both algorithms solve the unweighted version, they are not that closely related. Existing expositions convey intuition behind the choice of one or both of the orders, but do not explain where the difference arises and do not highlight the close relationship between the algorithms.
This note addresses those shortcomings. We derive the greedy algorithm from the dynamic program by simplifying the latter under the preconditions of the former. The change in order follows naturally. We apply a similar strategy to two other common examples, namely Knapsack, where we derive the greedy algorithm for the unit-value restriction from the dynamic program for the general case, and Shortest Paths, where we derive Dijkstra’s algorithm from the Bellman-Ford dynamic program.
@@
In a dynamic program, we pick a particular component, say $x_{1}$. We consider all possible settings $v_{1}$ for $y_{1}$, optimize $f_{x}$ with the additional constraint $y_{1}=v_{1}$, and assign $y_{1}$ to a value $v^{*}_{1}$ that produces the best result. The approach hinges on an efficient reduction from instances with the above additional constraint to proper $\Pi$-instances of size smaller than $n$. Repeated application yields a recursive algorithm for $\Pi$ that can be captured in the Bellman equations, which express the optimal value for input $x$ as the maximum or minimum of terms involving the optimal values for smaller instances. Correctness typically follows in a streamlined fashion. The name of the game is to ensure that the number of subproblems on any given input $x$ remains small, i.e., the number of distinct calls to $\Pi$ that arise throughout the entire recursive process on input $x$. This often requires reformulating the decisions to choose among, and sometimes also reordering the components. If and when the approach succeeds, memoization yields an efficient top-down algorithm for $\Pi$. Alternately, when the subproblems can be described in terms of $x$ and $k$ additional parameters, a $k$-dimensional table with the optimal values of the subproblems can be computed iteratively in a bottom-up fashion, after which a second phase retrieves decisions that achieve the optimal values in a top-down fashion.
@@
    Interval Scheduling. We start from a straightforward dynamic program for the weighted version and derive the classical greedy algorithm for the unweighed version. The weighted version represents a rare example of a computational problem where the order of the components in the dynamic program matters; earliest start time first yields the best bound on the number of subproblems. Our derivation of the greedy algorithm for the unweighted version transforms the order in a natural way into earliest finish time first. The monotonicity of the 1-dimensional table of subproblem solution values justifies the transition.
@@
    Knapsack. We start from the standard dynamic program for the general problem with integral item weights and arbitrary item values. Compared to Interval Scheduling, the subproblems have an additional parameter, namely, the weight limit of the knapsack. The order of the items plays no role in the analysis of the dynamic program. Our derivation for the case with unit values leads to lightest first as the greedy order. The justification follows from the monotonicity of the 2-dimensional table of subproblem solution values with respect to each parameter separately.
@@
The paradigms of dynamic programming and greed are studied in several areas, including algorithms and complexity, operations research and management science, and programming languages. The connections reported in this note ought to be known, but I have been unable to find them in the literature (including textbooks and lecture notes). The coverage in [CLRS22] gets closest in that it also relates dynamic programming and greed for Interval Scheduling and for Knapsack, but the perspective and connections are quite different from ours. For Interval Scheduling, [CLRS22] studies the unweighted version only, and only considers earliest finish time first, so there is no explanation for the switch in order. Their dynamic program tries every possible interval to reduce an instance into at most two smaller instances; they then argue that picking the first interval is always best. For Knapsack, [CLRS22] simplifies the problem by allowing fractional items rather than restricting to unit values, and explains why the fractional version allows a greedy approach whereas the original problem does not seem to. The presentation derives the greedy algorithm for the fractional variant from the dynamic program for the proper problem.
@@
For Shortest Paths, some connections between Bellman-Ford and Dijkstra are well-known, but our derivative approach seems new. The closest presentations [Den03, Ber17, Sni06] view Dijkstra as a modification of Bellman-Ford where the Bellman equations are applied best first (and each only once) instead of breadth first (and each possibly multiple times). They argue correctness of Dijkstra from scratch rather than deriving it from the correctness of Bellman-Ford.
@@
Much of the existing literature on the paradigms aims at precisely formalizing them. The programming languages community has developed very general formalisms within the framework of category theory to synthesize programs from problem specifications according to various algorithmic paradigms. A paper with a very similar title to ours [BdM93] builds dynamic programs and greedy algorithms under correctness conditions that take the form of monotonicity requirements in the categorical calculus of relations. The conditions for a greedy algorithm are stricter than for a dynamic program, but merely express the very existence of a local criterion that guarantees optimality, and the construction presupposes knowledge of the local criterion. As such, whereas the generality of the formalism is very suitable for program synthesis, the results provide little clue to an algorithm designer what greedy approaches might work or how to formulate a given computational problem to enable a greedy approach. Proofs of successful instantiations hinge on classical correctness arguments for known greedy algorithms. See Section 5 for more details.
@@
In a simpler framework for optimization problems with a decomposable structure, [Lew06] proposes a “canonical” way of deriving greedy algorithms from dynamic programs, but without correctness guarantees. The canonical local criterion is the straightforward one induced by the decomposition of the objective function. The framework is easy to apply but is much more limited than the categorical one and suffers from similar drawbacks. It provides no indication what order of components could work, and correctness proofs again boil down to classical ones.
@@
Although not the focus of this note, a potential research direction is to formally capture our approach as well as greedy-stays-ahead arguments, and to investigate their relationship in general. See also Section 5.
@@
The next three sections develop the three instantiations mentioned above. Each time we review the standard dynamic program for the general problem and derive from it the well-known greedy algorithm for the restricted problem. Readers who want to quickly get to the core, can skip the reviews and jump to Section 2.2, and then to Sections 3.2 and 4.2. The last section takes a closer look at the various monotonicity requirements involved, which relates to the open question.
@@ 2.1 Dynamic program
We consider component $x_{1}$, determine the optimum under the two choices for $y_{1}$, and take the best of the two. For either choice of $y_{1}$, the optimum value can be expressed in terms of the solution of an instance given by a subsequence of $x_{1},\dots,x_{n}$: the subsequence $x_{2},\dots,x_{n}$ for $y_{1}=0$, and the subsequence of the indices $j$ with $I_{1}\cap I_{j}=\emptyset$ for $y_{1}=1$. Repeated application of this reduction yields a recursive algorithm, which we can memoize. Since the order of the components within a sequence does not affect the outcome, the subproblems can all be described by subsets $J\subseteq[n]$. The crucial question is how many distinct subsets arise throughout the recursion.
@@
Figure 1: Ordering for Interval Scheduling instance yielding $2^{\Omega(n)}$ distinct subproblems.
At level $m$ of the recursion tree for the instance in Figure 1, every possible subset $J$ of $\{m+1,\dots,m\}$ arises as the remaining subproblem for a valid choice of decisions $y_{1},\dots,y_{m}$, namely the choice $y_{i}=0$ if $m+i\not\in J$, and $y_{i}=1$ otherwise. Thus, the number of distinct subproblems is at least $2^{m}=2^{n/2}$.
[…]
@@ 2.2 Greedy algorithm
[…]
itemSide effect:
We derive a greedy algorithm for the unit-value setting from the above dynamic program for the general setting. Consider the components ordered earliest start time first, as in the dynamic program. The recursive cases of the Bellman equations (2) can be simplified using the precondition $v\equiv 1$ and be rewritten as
@@
[…]
@@
Each term $1+\mathop{\mathrm{OPT}}(\mathrm{next}(j))$ in (itemSide effect:) corresponds to the situation where interval $I_{j}$ is the first interval in the given order that is selected from among $I_{i},\dots,I_{n}$. We can rewrite (itemSide effect:) as
@@
By definition (3), the value of $\mathrm{next}(j)$ is determined by the finish time $f_{j}$ of $I_{j}$, and is non-decreasing as a function of $f_{j}$ since increasing $f_{j}$ shrinks the set of components $k$ of which we take the minimum. Thus, $i^{*}$ is the (first) index of an interval $I_{j}$ with the earliest finish time among $I_{i},\dots,I_{n}$. The interpretation of the terms in (itemSide effect:) then tells us that an optimal solution is to select interval $I_{i^{*}}$ and then continue with the remaining intervals that do not overlap with $I_{i^{*}}$. That is to say, we greedily consider the intervals earliest finish time first, which is exactly what the classical greedy algorithm does.
@@ 2.3 Order considerations
We end our discussion of Interval Scheduling with a few additional remarks about the different orders involved. The order of components that we mention is always the one for applying the reduction that underlies both paradigms. In the dynamic programming, this order is used in the memoized process and in the second phase (solution retrieval) of the iterative implementation (see Algorithm 2). The first phase of the iterative implementation considers the components in the reverse order (see Algorithm 1).
The reverse order of the dynamic program for Interval Scheduling is earliest start time last. Upon reversal of the time line, this becomes earliest finish time first. The orientation of the time line does affect correctness and an increasing loop variable is more natural than a decreasing one. As such, rather than sorting components earliest deadline first as in Algorithm 1, typical implementations of the first phase of the dynamic program for Interval Scheduling sort the components earliest finish time first, which is the greedy order for Unit-Value Interval Scheduling. This agreement is a coincidence. When teaching greed before dynamic programming, picking the greedy order for Unit-Value Interval Scheduling as the order in the first iterative phase of the dynamic program for Interval Scheduling requires justification. The justification for the greedy order is correctness, whereas the justification for the order in the dynamic program is efficiency.
@@
[…]
Figure 2: Interval Scheduling instance where earliest start time first yields $\Theta(n^{2})$ distinct subproblems.
The instances in Figure 2 have even $n=2m$, intervals ordered earliest finish time first, and unspecified values. For each $i\in[m]$, all subproblems with parameters $(\ell,i)$ for $0\leq\ell\leq i-1$ are present at level $i-1$ of the recursion tree. They correspond to subsets of the form $J=\{i,\dots,m\}\cup\{m+\ell+1,\dots,2m\}$ and occur when we set $y_{\ell}=1$ (for $\ell\geq 1$) and $y_{\ell+1}=\dots=y_{i-1}=0$. There are $\binom{m+1}{2}=\Theta(n^{2})$ such subproblems in total.
@@ 3 Knapsack
itemSide effect:
Similar to Interval Scheduling, the input can be viewed as consisting of the components $x_{i}=(w_{i},v_{i})$ for $i\in[n]$ and the additional entity $W$. We need to make a binary decision $y_{i}$ about each component $x_{i}$, where $y_{i}$ indicates whether item $i$ is included in the knapsack. Alternately, the sequence of binary decisions can be replaced by successive non-binary decisions as to which item to include next (if any). We again adopt the former view.
@@ 3.1 Dynamic program
Given a decision for the first item, it remains to solve the instance consisting of items 2 through $n$ with weight limit $W$ in case we do not include item 1, and the same instance but with a reduced weight limit of $W-w_{1}$ in case we include item 1 (which is only an option for $w_{1}\leq W$). The subproblems that arise through repeated application of the reduction consist of a suffix of the given list of items, and a nonnegative integer at most $W$ as the weight limit. The solutions can be represented in a matrix $\mathop{\mathrm{OPT}}$. For $i\in[n+1]$ and $w\in\{0,\dots,W\}$
$\mathop{\mathrm{OPT}}(i,w)\doteq\text{optimal value for components }x_{i},\dots,x_{n}\text{ and weight limit }w.$ (10)

### v2

@@ 1 Introduction
Many an undergraduate algorithms course features Interval Scheduling to illustrate the design paradigms of greed and dynamic programming. The unweighted version admits an appealing greedy strategy that can be shown correct via several types of arguments (charging, exchange, greedy stays ahead). The weighted version exhibits a straightforward principle of optimality and demonstrates one of the measures that can be taken to keep the number of subinstances in a dynamic program small – ordering the input components judiciously. However, the appropriate order for the memoized dynamic program differs from the order in the greedy algorithm. This tends to confuse students and gives some the impression that, even though both algorithms solve the unweighted version, they are not that closely related. Existing expositions convey intuition behind the choice of one or both of the orders, but do not explain where the difference arises and do not highlight the close relationship between the algorithms.
This paper addresses those shortcomings. We derive the greedy algorithm from the dynamic program by simplifying the latter under the preconditions of the former. The change in order follows naturally. We apply a similar strategy to two other common examples, namely Knapsack, where we derive the greedy algorithm for the unit-value restriction from the dynamic program for the general case, and Shortest Paths, where we derive Dijkstra’s algorithm from the Bellman-Ford dynamic program.
@@
In a dynamic program, we pick a particular component, say $x_{1}$. We consider all possible settings $v_{1}$ for $y_{1}$, optimize $f_{x}$ with the additional constraint $y_{1}=v_{1}$, and assign $y_{1}$ to a value $v^{*}_{1}$ that produces the best result. The approach hinges on an efficient reduction from instances with the above additional constraint to proper $\Pi$-instances of size smaller than $n$. Repeated application yields a recursive algorithm for $\Pi$ that can be captured in the Bellman equations, which express the optimal value for input $x$ as the maximum or minimum of terms involving the optimal values for smaller instances. Correctness typically follows in a streamlined fashion. The name of the game is to ensure that the number of subinstances on any given input $x$ remains small, i.e., the number of distinct calls to $\Pi$ that arise throughout the entire recursive process on input $x$. This often requires reformulating the decisions to choose among, and sometimes also reordering the components. If and when the approach succeeds, memoization yields an efficient top-down algorithm for $\Pi$. Alternately, when the subinstances can be described in terms of $x$ and $k$ additional parameters, a $k$-dimensional table with the optimal values of the subinstances can be computed iteratively in a bottom-up fashion, after which a second phase retrieves decisions that achieve the optimal values in a top-down fashion.
@@
    Interval Scheduling. We start from a straightforward dynamic program for the weighted version and derive the classical greedy algorithm for the unweighed version. The weighted version represents a rare example of a computational problem where the order of the components in the dynamic program matters; earliest start time first yields the best bound on the number of subinstances. Our derivation of the greedy algorithm for the unweighted version transforms the order in a natural way into earliest finish time first. The monotonicity of the 1-dimensional table of subinstance solution values justifies the transition.
@@
    Knapsack. We start from the standard dynamic program for the general problem with integral item weights and arbitrary item values. Compared to Interval Scheduling, the subinstances have an additional parameter, namely, the weight limit of the knapsack. The order of the items plays no role in the analysis of the dynamic program. Our derivation for the case with unit values leads to lightest first as the greedy order. The justification follows from the monotonicity of the 2-dimensional table of subinstance solution values with respect to each parameter separately.
@@
The paradigms of dynamic programming and greed are studied in several areas, including algorithms and complexity, operations research and management science, and programming languages. The connections reported in this paper ought to be known, but I have been unable to find them in the literature (including textbooks and lecture notes). The coverage in [CLRS22] gets closest in that it also relates dynamic programming and greed for Interval Scheduling and for Knapsack, but the perspective and connections are quite different from the approach here. For Interval Scheduling, [CLRS22] studies the unweighted version only, and only considers earliest finish time first, so there is no explanation for the switch in order. Their dynamic program tries every possible interval to reduce an instance into at most two smaller instances; they then argue that picking the first interval is always best. For Knapsack, [CLRS22] simplifies the problem by allowing fractional items rather than restricting to unit values, and explains why the fractional version allows a greedy approach whereas the original problem does not seem to. The presentation derives the greedy algorithm for the fractional variant from the dynamic program for the proper problem.
@@
For Shortest Paths, some connections between Bellman-Ford and Dijkstra are well-known, but the derivative approach in this paper seems new. The closest presentations [Den03, Ber17, Sni06] view Dijkstra as a modification of Bellman-Ford where the Bellman equations are applied best first (and each only once) instead of breadth first (and each possibly multiple times). They argue correctness of Dijkstra from scratch rather than deriving it from the correctness of Bellman-Ford.
We note that there are ways in which Dijkstra’s algorithm itself resembles a dynamic program. In the single-pair version, whether to include an edge on the shortest path from $s$ to $t$ is not decided as soon as the edge is considered, but in a second phase whose order is the reverse of the greedy order. Arguably, the single-source version that we consider is greedy. In every step, a shortest path to some new target $t$ is completed by the choice of a particular final edge.
@@
Much of the existing literature on the paradigms aims at precisely formalizing them. The programming languages community has developed very general formalisms within the framework of category theory to synthesize programs from problem specifications according to various algorithmic paradigms. A paper with a very similar title to this work builds dynamic programs and greedy algorithms under correctness conditions that take the form of monotonicity requirements in the categorical calculus of relations [BdM93]. The conditions for a greedy algorithm are stricter than for a dynamic program, but merely express the very existence of a local criterion that guarantees optimality, and the construction presupposes knowledge of the local criterion. As such, whereas the generality of the formalism is suitable for program synthesis, the results provide little clue to an algorithm designer what greedy approaches might work or how to formulate a given computational problem to enable a greedy approach. Proofs of successful instantiations hinge on classical correctness arguments for known greedy algorithms. See Section 5 for more details.
@@
In a simpler framework for optimization problems with a decomposable structure, [Lew06] proposes a “canonical” way of deriving greedy algorithms from dynamic programs, but without correctness guarantees. The canonical local criterion is the straightforward one induced by the decomposition of the objective function. The framework is easy to apply but is more limited than the categorical one and suffers from similar drawbacks. It provides no indication what order of components could work, and correctness proofs again boil down to classical ones.
@@
A potential research direction is to formally capture the approach here as well as greedy-stays-ahead arguments, and to investigate their relationship in general. See also Section 5.
@@
The next three sections develop the three instantiations mentioned above. Each time we review the standard dynamic program for the general problem and derive from it the well-known greedy algorithm for the restricted problem. Readers who want to quickly get to the core, can skip the reviews and jump to Section 2.2, and then to Sections 3.2 and 4.2. Section 5 takes a closer look at the various monotonicity requirements involved, which relates to the open question.
@@ 2.1 Dynamic program
We consider component $x_{1}$, determine the optimum under the two choices for $y_{1}$, and take the best of the two. For either choice of $y_{1}$, the optimum value can be expressed in terms of the solution of an instance given by a subsequence of $x_{1},\dots,x_{n}$: the subsequence $x_{2},\dots,x_{n}$ for $y_{1}=0$, and the subsequence of the indices $j$ with $I_{1}\cap I_{j}=\emptyset$ for $y_{1}=1$. Repeated application of this reduction yields a recursive algorithm, which we can memoize. Since the order of the components within a sequence does not affect the outcome, the subinstances can all be described by subsets $J\subseteq[n]$. The crucial question is how many distinct subsets arise throughout the recursion.
@@
Figure 1: Ordering for Interval Scheduling instance yielding $2^{\Omega(n)}$ distinct subinstances.
At level $m$ of the recursion tree for the instance in Figure 1, every possible subset $J$ of $\{m+1,\dots,2m\}$ arises as the remaining subinstance for a valid choice of decisions $y_{1},\dots,y_{m}$, namely the choice $y_{i}=0$ if $m+i\not\in J$, and $y_{i}=1$ otherwise. Thus, the number of distinct subinstances is at least $2^{m}=2^{n/2}$.
However, when we order the components earliest start time first, every subset $J$ in the recursion tree has the structure of a suffix of $[n]$, of which there are only $n+1$. This is because when we apply the reduction to a subinstance corresponding to the suffix $J\doteq\{i,\dots,n\}$ (initially $i=1$), either choice of $y_{i}$ merely removes a prefix from the suffix $J$: the prefix of length 1 for $y_{i}=0$, and the prefix consisting of all intervals with start time before $f_{i}$ for $y_{i}=1$. This leads to a 1-dimensional table $\mathop{\mathrm{OPT}}$ of subinstance solutions: For $i\in[n+1]$
@@ 2.2 Greedy algorithm
$S\subseteq[n]$ such that $|S|$ is maximized and $(\forall i\neq j\in S)\,I_{i}\cap I_{j}=\emptyset$
We derive a greedy algorithm for the unit-value setting from the above dynamic program for the general setting. Consider the components ordered earliest start time first, as in the dynamic program. The recursive cases of the Bellman equations (2) can be simplified using the precondition $v\equiv 1$ and rewritten as
@@
$\mathop{\mathrm{OPT}}(i)$ $=$ $\max\Big(\,1+\mathop{\mathrm{OPT}}(\mathrm{next}(i)),$
@@
Each term $1+\mathop{\mathrm{OPT}}(\mathrm{next}(j))$ in (itemOutput:) corresponds to the situation where interval $I_{j}$ is the first interval in the given order that is selected from among $I_{i},\dots,I_{n}$. We can rewrite (itemOutput:) as
@@
By definition (3), the value of $\mathrm{next}(j)$ is determined by the finish time $f_{j}$ of $I_{j}$, and is non-decreasing as a function of $f_{j}$ since increasing $f_{j}$ shrinks the set of components $k$ of which we take the minimum. Thus, $i^{*}$ is the (first) index of an interval $I_{j}$ with the earliest finish time among $I_{i},\dots,I_{n}$. The interpretation of the terms in (itemOutput:) then tells us that an optimal solution is to select interval $I_{i^{*}}$ and then continue with the remaining intervals that do not overlap with $I_{i^{*}}$. That is to say, we greedily consider the intervals earliest finish time first, which is exactly what the classical greedy algorithm does.
@@ 2.3 Order considerations
We end our discussion of Interval Scheduling with a few additional remarks about the different orders involved. The order of components that we mention is the one for applying the reduction that underlies both paradigms. In the dynamic program, this order is used in the memoized process and in the second phase (solution retrieval) of the iterative implementation (see Algorithm 2). The first phase of the iterative implementation considers the components in the reverse order (see Algorithm 1).
The reverse order in the dynamic program for Interval Scheduling is earliest start time last. Upon reversal of the time line, this becomes earliest finish time first. The orientation of the time line does not affect correctness. As an increasing loop variable is more natural than a decreasing one, rather than sorting components earliest deadline first as in Algorithm 1, typical implementations of the first phase of the dynamic program for Interval Scheduling sort the components earliest finish time first. This agrees with the greedy order for Unit-Value Interval Scheduling, but the agreement is a coincidence. When teaching greed before dynamic programming, picking the greedy order for Unit-Value Interval Scheduling as the order in the first iterative phase of the dynamic program for Interval Scheduling requires justification. The justification for the greedy order is correctness, whereas the justification for the order in the dynamic program is efficiency.
To complete the picture, we analyze ordering the components earliest finish time first in the dynamic program for Interval Scheduling. The subsets $J$ are no longer merely suffixes of $[n]$, but still have enough structure to keep the number of subinstances small, albeit larger than for earliest start time first. Each $J$ now has the form $\{j\in\{i,\dots,n\}:s_{j}\geq f_{\ell}\}$ for $0\leq\ell<i\leq n$, where $f_{0}\doteq$-$\infty$ acts as a sentinel. The parameter space is therefore 2-dimensional and of size $O(n^{2})$. Figure 2 describes instances for which $\Omega(n^{2})$ of the combinations $(\ell,i)$ actually occur. This means that the number of subinstances for earliest finish time first is $\Theta(n^{2})$ as opposed to $\Theta(n)$ for earliest start time first, resulting in a running time of $\Theta(n^{2})$ in lieu of $\Theta(n\log n)$.
@@
Figure 2: Interval Scheduling instance where earliest start time first yields $\Theta(n^{2})$ distinct subinstances.
The instances in Figure 2 have even $n=2m$, intervals ordered earliest finish time first, and unspecified values. For each $i\in[m]$, all subinstances with parameters $(\ell,i)$ for $0\leq\ell\leq i-1$ are present at level i1i-1 of the recursion tree. They correspond to subsets of the form $J=\{i,\dots,m\}\cup\{m+\ell+1,\dots,2m\}$ and occur when we set $y_{\ell}=1$ (for $\ell\geq 1$) and $y_{\ell+1}=\dots=y_{i-1}=0$. There are $\binom{m+1}{2}=\Theta(n^{2})$ such subinstances in total.
@@ 3 Knapsack
Similar to Interval Scheduling, the input can be viewed as consisting of the components $x_{i}=(w_{i},v_{i})$ for $i\in[n]$ and the additional quantity $W$. We need to make a binary decision $y_{i}$ about each component $x_{i}$, where $y_{i}$ indicates whether item $i$ is included in the knapsack. Alternately, the sequence of binary decisions can be replaced by successive non-binary decisions as to which item to include next (if any). We again adopt the former view.
@@ 3.1 Dynamic program
Given a decision for the first item, it remains to solve the instance consisting of items 2 through $n$ with weight limit $W$ in case we do not include item 1, and the same instance but with a reduced weight limit of $W-w_{1}$ in case we include item 1 (which is only an option if $w_{1}\leq W$). The subinstances that arise through repeated application of the reduction consist of a suffix of the given list of items, and a nonnegative integer at most $W$ as the weight limit. The solutions can be represented in a matrix $\mathop{\mathrm{OPT}}$. For $i\in[n+1]$ and $w\in\{0,\dots,W\}$
[…]